# LUMOS: Large User MOdel Series for User Behavior Prediction

### v1

@@ 5 Results and Analysis
To validate whether offline performance improvements translate to measurable business impact, we deployed LUMOS predictions in production via an intelligent offers strategy. A controlled 18-day A/B test with approximately 700,000 users per cohort demonstrated that LUMOS-powered offers led to a 3.15% increase in Daily Active Users (DAU) while simultaneously reducing offer expenditure by 2.47%, confirming that offline gains translate directly to enhanced user engagement and operational efficiency.
5.1 Attention Pattern Analysis
@@
Figure 4: Cross-attention weights when predicting user behaviour at the start of IPL season.
Figure 4 reveals three distinct learned patterns: (1) Recency focus with strong attention to the last 7-14 days (positions 346-360), (2) Event-specific memory showing peaks at positions corresponding to previous year’s IPL final (position 42), T20 World Cup final (position 156), and start of last IPL season (position 298), and (3) Seasonal patterns with moderate attention to the same calendar period from the previous year. This demonstrates that the model learns meaningful temporal dependencies without explicit programming, automatically identifying that past tournament behaviour predicts future tournament engagement.
5.2 Scaling Laws Analysis
To understand LUMOS’s scaling behavior and validate its architecture against established neural network principles, we analyze how model performance scales with both training data amount and model size as was pioneered by Hoffmann et al. [2]. Using a fixed 7.23M parameter model across 149 well-trained runs with data ranging from 2,532 GB to 50,000 GB, we observe a power law relationship $L(D)=0.507\times D^{-0.048}$ with scaling exponent $\alpha=-0.048$ ($R^{2}=0.869$, $p<0.001$), indicating diminishing returns with increasing data—doubling data yields approximately 3.3% loss reduction. For model capacity, examining 14 model sizes from 7.23M to 20.81M parameters reveals a stronger scaling relationship $L(N)=0.438\times N^{-0.152}$ with exponent $\alpha=-0.152$ ($R^{2}=0.622$). The model size scaling exponent is approximately 3× stronger than the data scaling exponent, demonstrating that architectural capacity has greater impact on performance than data quantity in the observed regime.
@@ 6 Conclusion
The key technical contributions of LUMOS include: (1) a novel cross-attention mechanism that conditions predictions on future known events, enabling the model to learn complex temporal relationships between user history and upcoming event context; (2) a multi-modal tokenization scheme that unifies user transactions, event context, and static features into rich daily representations; and (3) a production-ready implementation at scale - all of which are backed by empirical evidence - both offline and online.
@@
Our results reveal several important auxiliary insights. First, the bidirectional importance of event context — where both past and future event context play a crucial role in predicting user behaviour. Second, learned positional embeddings significantly outperform generic alternatives, indicating that user behaviour exhibits unique temporal patterns requiring task-specific modeling. Third, our scaling laws analysis demonstrates that LUMOS follows predictable power law relationships, with model size scaling ($\alpha=-0.152$) providing 3× stronger improvements than data scaling ($\alpha=-0.048$), validating the architectural approach and providing clear guidance for future scaling decisions.
@@
As user behaviour data continues to grow in volume and complexity, approaches like LUMOS that can automatically discover patterns from raw data will become vital. By demonstrating that transformers can successfully model user behaviour at scale, we hope to inspire further research at the intersection of deep learning and user behaviour modelling.

### v2

@@ 5 Results and Analysis
To validate whether offline performance improvements translate to measurable business impact, we deployed LUMOS predictions in production via Pacman—our proprietary intelligent offers strategy. A controlled 18-day A/B test with approximately 700,000 users per cohort demonstrated that LUMOS-powered offers led to a 3.15% increase in Daily Active Users (DAU) while simultaneously reducing offer expenditure by 2.47%, confirming that offline gains translate directly to enhanced user engagement and operational efficiency.
5.1 Attention Patterns
@@
Figure 4: Cross-attention weights when predicting user behaviour at the start of IPL season. The heatmap reveals three distinct learned patterns: (1) Recency focus with strong attention to the last 7-14 days (positions 346-360), (2) Event-specific memory showing peaks at positions corresponding to previous year’s IPL final (position 42), T20 World Cup final (position 156), and start of last IPL season (position 298), and (3) Seasonal patterns with moderate attention to the same calendar period from the previous year. This demonstrates that the model learns meaningful temporal dependencies without explicit programming, automatically identifying that past tournament behaviour predicts future tournament engagement.
Figure 4 reveals how the model automatically learns to attend to relevant historical patterns.
5.2 Scaling Laws
To understand LUMOS’s scaling behavior and validate its architecture against established neural network principles, we analyze how model performance scales with both training data amount and model size as was pioneered by the Chinchilla paper [2]. Using a fixed 7.23M parameter model across 9 well-trained runs with data ranging from 2,532 GB to 50,000 GB, we observe a power law relationship $L(D)=0.507\times D^{-0.048}$ with scaling exponent $\alpha=-0.048$ ($R^{2}=0.869$, $p<0.001$), indicating diminishing returns with increasing data—doubling data yields approximately 3.3% loss reduction.
For model capacity, examining 4 model sizes from 7.23M to 20.81M parameters reveals a stronger scaling relationship $L(N)=0.438\times N^{-0.152}$ with exponent $\alpha=-0.152$ ($R^{2}=0.622$). The model size scaling exponent is approximately 3× stronger than the data scaling exponent, demonstrating that architectural capacity has greater impact on performance than data quantity in the observed regime.
Figure 5: Neural network scaling laws for LUMOS: (a) Loss vs. training data amount, (b) Loss vs. model size. Both exhibit clear power law relationships.
Table 2: Model configurations tested for scaling analysis
Config Parameters $d_{model}$ Layers (E/D) Heads Val Loss
Small 7.23M 512 6/6 8 0.3347
Medium 12.18M 640 8/8 8 0.3298
Large 16.42M 768 8/8 12 0.3271
X-Large 20.81M 768 10/10 12 0.3254
The Chinchilla scaling laws [2] found exponents of $\alpha\approx-0.34$ for data and $\alpha\approx-0.28$ for model size in language models. LUMOS exhibits weaker scaling ($\alpha_{data}=-0.048$, $\alpha_{model}=-0.152$), suggesting user behavior prediction may be more sample-efficient than language modeling but still benefits from architectural capacity.
@@ 6 Conclusion
The key technical contributions of LUMOS include: (1) a novel cross-attention mechanism that conditions predictions on future known events, enabling the model to learn complex temporal relationships between user history and upcoming event context; (2) a multi-modal tokenization scheme that unifies user activities, event context, and static features into rich daily representations; and (3) a production-ready implementation at scale - all of which are backed by empirical evidence - both offline and online.
@@
Our results reveal several important auxiliary insights. First, the bidirectional importance of event context — where both past and future event context play a crucial role in predicting user behaviour. Second, learned positional embeddings significantly outperform generic alternatives, indicating that user behaviour exhibits unique temporal patterns requiring task-specific modeling. Third, our scaling laws demonstrate that LUMOS follows predictable power law relationships, with model size scaling ($\alpha=-0.152$) providing 3× stronger improvements than data scaling ($\alpha=-0.048$), validating the architectural approach and providing clear guidance for future scaling decisions.
@@
As user behaviour data continues to grow in volume and complexity, approaches like LUMOS that can automatically discover patterns from raw data will become increasingly vital. By demonstrating that transformers can successfully model user behaviour at scale, we hope to inspire further research at the intersection of deep learning and user analytics. The shift from manual feature engineering to learned representations is not just a technical improvement—it fundamentally changes how we build and deploy user understanding systems, enabling more accurate, adaptable, and maintainable solutions for the challenges of modern digital platforms.
Appendix 0.A Positional Embedding Analysis
0.A.1 Frequency Components Analysis
To understand what temporal patterns the learned positional embeddings capture, we apply Fast Fourier Transform (FFT) to extract frequency components from both learned and sinusoidal embeddings.
Figure 6: Frequency spectrum analysis: (Left) Sinusoidal embeddings distribute energy across all frequencies. (Right) Learned embeddings concentrate in low frequencies (<< 0.1 cycles/day), suggesting the model prioritizes broad temporal context over precise positioning.
0.A.2 Key Findings
Learned embeddings concentrate >>80% of their energy in frequencies below 0.1 cycles/day (periods >>10 days). High-frequency components (>>0.5 cycles/day) are nearly absent.This frequency distribution suggests that for user behavior prediction, distinguishing ”last week” from ”last month” is more important than distinguishing ”day 42” from ”day 45”. The model learns to encode coarse temporal granularity rather than precise day-level positioning. Fixed sinusoidal embeddings distribute energy uniformly across all frequencies, potentially encoding unnecessary high-frequency information. The learned embeddings adaptively focus on task-relevant temporal scales.
0.A.3 Comparison of Positional Encoding Strategies
We systematically evaluated five positional encoding strategies on the user behavior prediction task:
Table 3: Performance comparison of positional encoding strategies
Strategy Validation Loss Description
None 0.3412 No positional information
Sinusoidal 0.3389 Fixed trigonometric functions
RoPE 0.3365 Rotary position embeddings
TAPE 0.3358 Time-aware position embeddings
Learned 0.3347 Trainable embeddings (used in LUMOS)
Appendix 0.B Embedding Extraction and Analysis
Embedding Quality: We extracted user embeddings from the trained multi-task model according to the method described in Section 0.B.1 and evaluated their predictive performance on downstream tasks to assess generalizability.
0.B.1 Extracting User Embeddings
Beyond predictions, LUMOS generates valuable user embeddings for downstream tasks. We extract representations at multiple levels:
  • Static embedding: $\mathbf{h}^{static}$ for cold-start scenarios with minimal user information
  • Dynamic User embedding: Aggregated encoder output capturing current temporal user state
  • Event context embedding: Learned representation of event context
0.B.1.1 Dynamic User Embeddings
A key innovation in LUMOS is the generation of dynamic user embeddings that are refreshed daily. These embeddings capture the evolving state of each user based on their most recent behavioral patterns:
$\mathbf{e}^{user}_{t}=\text{Aggregate}(\mathbf{H}^{enc}_{t})$ (11)
where $\mathbf{H}^{enc}_{t}$ is the encoder output at time $t$ incorporating all historical information up to that point.
For aggregation, we primarily use exponentially weighted averaging:
$\mathbf{e}^{user}_{t}=\sum_{i=1}^{T_{hist}}\alpha^{T_{hist}-i}\mathbf{H}^{enc}_{t,i}$ (12)
where $\alpha\in(0,1)$ is a decay factor that emphasizes recent history. This approach ensures that the embedding reflects the user’s current state while retaining important historical patterns.
These dynamic embeddings serve multiple purposes, including identifying cohorts with similar behavioral patterns through user similarity detection, flagging sudden changes in user behavior via anomaly detection, tailoring content and offers based on the current user state through personalization, and providing rich inputs for downstream ML models via feature extraction.
0.B.2 User Embedding Analysis
0.B.2.1 Embedding Quality on Downstream Tasks
We evaluate learned user embeddings on five downstream tasks using simple logistic regression:
Table 4: User embedding performance on downstream tasks (ROC-AUC)
Embedding Source Task 1 Task 2 Task 3 Task 4 Task 5
Single-task (Task 1 only) 0.906 0.654 0.411 0.657 0.974
Single-task (Task 2 only) 0.879 0.672 0.425 0.641 0.961
Multi-task LUMOS 0.977 0.722 0.442 0.708 0.975
Relative Improvement +7.8% +7.4% +4.0% +7.8% +0.1%
Multi-task embeddings dramatically outperform single-task specialists, even on the tasks the specialists were trained for. This suggests that joint training forces the model to learn more generalizable representations that capture fundamental aspects of user behavior rather than task-specific patterns. We found that exponentially weighted aggregation with high decay factor performs best, confirming that recent behavior is most predictive while historical context provides important background.
0.B.3 Event Context Embedding Analysis
The learned event context embeddings ($d_{supply\_embed}=64$) capture meaningful event context. Each day’s event context features (match counts, tournament indicators) are embedded into a continuous vector space where semantically similar days cluster together.
Figure 7: t-SNE projection of daily event context embeddings reveals distinct clusters for days with different IPL match counts: (blue) no IPL matches, (orange) single IPL match, (green) IPL double-headers. The learned space captures event context similarity without explicit supervision.
Use Cases for Event Context Embeddings:
  • Event Context Similarity Analysis: Find historical days with similar event calendars
  • What-If Scenarios: Embed hypothetical future event context to simulate user responses
  • Event Impact Assessment: Measure distance between normal days and tournament days
  • Communications Targeting: Identify users most responsive to specific event context patterns
0.B.4 Ablation Study: Impact of Event Context Data
We investigate the importance of event context by systematically removing past and future event context information:
Table 5: Event context ablation results on churn prediction
Configuration Validation Loss ROC-AUC $\Delta$ Loss
Full Model (Past + Future Event Context) 0.3331 0.9300 -
No Future Event Context 0.3350 0.9294 +0.57%
No Past Event Context 0.3349 0.9297 +0.54%
No Event Context (Past or Future) 0.3345 0.9298 +0.42%
Surprisingly, removing either past or future event context individually hurts performance more than removing both. This suggests that past and future event context provide complementary information—the model learns to associate historical user responses to past events with likely responses to similar future events. When only one is available, the model cannot make these connections effectively.
Appendix 0.C Production Inference Infrastructure
0.C.1 Batch Inference Architecture
LUMOS operates in daily batch inference mode, scoring all active users ($\sim$5M/day) each morning. The inference pipeline consists of five stages:
[Inference Pipeline Diagram]
Data Fetch $\rightarrow$ Feature Scaling $\rightarrow$ Batch Inference $\rightarrow$ Post-processing $\rightarrow$ Storage
Figure 8: Production inference pipeline: Data fetch $\rightarrow$ Feature scaling $\rightarrow$ Batch inference $\rightarrow$ Post-processing $\rightarrow$ Storage.
Appendix 0.D Implementation Details